# CoRe: Context-Robust Remasking for Diffusion Language Models

### v1

@@ Abstract
Standard decoding in Masked Diffusion Models (MDMs) is hindered by context rigidity: tokens are retained based on transient high confidence, often ignoring that early predictions lack full context. This creates cascade effects where initial inconsistencies misguide the remaining generation. Existing revision strategies attempt to mitigate this by relying on static confidence scores, but these signals are inherently myopic; inconsistent tokens frequently appear confident to the model itself. To address this, we propose Context-Robust Remasking (CoRe), a training-free framework for inference-time revision. Rather than trusting static token probabilities, we identify context-brittle tokens by probing their sensitivity to targeted perturbations. We formalize revision as a robust optimization problem targeting worst-case context shifts. CoRe efficiently approximates this objective to expose unstable tokens, prioritizing them for revision. On LLaDA-8B-Base, CoRe delivers consistent improvements across reasoning and code benchmarks, outperforming compute-matched baselines and performance on code generation (MBPP) by up to $9.2\%$.
@@ 3 Problem Formulation
Context Rigidity Anchors Structural Inconsistencies.
Standard sampling treats each unmasked token as a permanent constraint: once a position is finalized, it is typically not revisited. This creates context rigidity. Token predictions at early steps are made when the context is sparse and ambiguity is high, yet they become fixed anchors for all subsequent steps. If an early generated token is suboptimal, later tokens are generated to remain consistent in context, leading to self-reinforcing inconsistencies that additional forward steps cannot rectify.
Uncertainty Heuristics Become Stale.
A natural remedy for context rigidity is inference-time revision: selectively resetting previously generated positions to [MASK] so the model can re-evaluate them. The central difficulty lies in identifying which tokens to revise. Existing strategies often repurpose stepwise heuristics such as low confidence [25] or top-$k$ margin [17] to select targets for remasking [35]. However, these signals quickly become stale: they reflect the ambiguity present when a token was originally sampled, failing to distinguish between tokens that are genuinely inconsistent with the updated context and those that were simply uncertain during intermediate steps.
@@
We propose to select revisions by measuring the sensitivity to context change. A generated token is considered stable if it remains strongly predicted even when parts of the surrounding context are masked; a token is brittle if its probability collapses under these perturbations, revealing that it is not robustly anchored. This shifts the selection criterion from “Was the token uncertain when it was chosen?” to “Does the token remain plausible under dynamic context change?” Our goal is to identify these brittle tokens under a fixed budget, enabling the decoder to revise tokens that are most context-sensitive across different decoding steps as the context evolves.

### v2

@@ Abstract
Standard decoding in Masked Diffusion Models (MDMs) is hindered by context rigidity: tokens are retained based on transient high confidence, often ignoring that early predictions lack full context. This creates cascade effects where initial inconsistencies misguide the remaining generation. Existing revision strategies attempt to mitigate this by relying on static confidence scores, but these signals are inherently short-sighted; inconsistent tokens frequently appear confident to the model itself. To address this, we propose Context-Robust Remasking (CoRe), a training-free framework for inference-time revision. Rather than trusting static token probabilities, we identify context-brittle tokens by probing their sensitivity to targeted perturbations. We formalize revision as a robust optimization problem targeting worst-case context shifts. CoRe efficiently approximates this objective to expose unstable tokens, prioritizing them for revision. On LLaDA-8B-Base, CoRe delivers consistent improvements across reasoning and code benchmarks, outperforming compute-matched baselines and boosting performance on code generation (MBPP) by up to $9.2\%$.
@@ 3 Problem Formulation
Context Rigidity Makes Revision Target Selection Hard.
Standard sampling treats unmasked tokens as immutable constraints; once finalized, a position is rarely revisited. This induces context rigidity: early predictions, conditioned on sparse context, become anchors for later steps. Consequently, a suboptimal early token forces subsequent generation to conform, yielding self-reinforcing inconsistencies. A natural remedy is inference-time revision—selectively resetting unmasked positions to [MASK]—yet the main challenge lies in identifying which tokens to revisit. Existing training-free strategies often rely on token-local uncertainty proxies such as low confidence [25, 35] or small top-$k$ margins [17]. However, these signals become stale: they reflect transient ambiguity at the moment of selection, often failing to detect tokens that become incompatible as the surrounding context evolves.
@@
Instead, we propose to select revisions by measuring the sensitivity to context change. A generated token is considered stable if it remains strongly predicted even when parts of the surrounding context are masked; a token is brittle if its probability collapses under these perturbations, revealing that it is not robustly anchored. This shifts the selection criterion from “Was the token uncertain when it was chosen?” to “Does the token remain plausible under dynamic context change?” Our goal is to identify these brittle tokens under a fixed budget, enabling the decoder to revise tokens that are most context-sensitive across different decoding steps as the context evolves.